Las Paletas Corporation has two different bonds currently outstanding. Bond M has a face value of $20,000 and matures in 20 years. The bond makes no payments for the first six years, then pays $1,100 every six months over the subsequent eight years, and finally pays $1,400 every six months over the last six years. Bond N also has a face value of $20,000 and a maturity of 20 years; it makes no coupon payments over the life of the bond. The required return on both these bonds is 6 percent compounded semiannually.

What is the current price of bond M and bond N?
Answer #1

Rate semiannual =6%/2 =3%
Price of Bond M =PV of Coupons of 1100 +PV of Coupons of 1400 + PV of Par Value
=1100*((1-(1+3%)^-16)/(3%*(1+3%)^12)+1400*(((1-(1+3%)^-12)/(3%*(1+3%)^28)+20000/(1+3%)^40 =21913.18

Price of Bond N =20000/(1+3%)^40 =6131.14